OpenStudy (anonymous):

Rationalize the denominator in the following expression. Use sqrt(x) for \[\sqrt{x}\] and root (x)(y) for \[\sqrt[x]{y}\]. 4/cube root (9)

OpenStudy (anonymous):

\[4/ \sqrt[3]{9}\]

OpenStudy (anonymous):

multiply top and bottom by \[\sqrt[3]{3}\]

OpenStudy (anonymous):

get \[\frac{4\sqrt[3]{x}}{\sqrt[3]{9}\times \sqrt[3]{3}}\] \[=\frac{4\sqrt[3]{3}}{\sqrt[3]{27}}=\frac{4\sqrt[3]{3}}{3}\]